OpenStudy (anonymous):

Point R divides EF¯¯¯¯¯ in the ratio 1 : 5. If the coordinates of E and F are (4, 8) and (11, 4), respectively, what are the coordinates of R to two decimal places?

OpenStudy (campbell_st):

well the 1st question is wether the division is internal, between EF, or external, so R is beyond E or F. so for internal division I start by writing equivalent ratios \[\frac{ER}{RF} = \frac{1}{5}\] then cross multiply so \[5ER = 1RF\] so you need to multiply the E point by 5 and the F point by 1 then its \[x = \frac{5E + F}{ 1 + 5}~~~~y = \frac{5E + F}{1 + 5}\] so multiply the E point by 5, both x and y then treat x and y in the same way using the formula... Hope it helps

OpenStudy (campbell_st):

you always divide by the sum of the parts in the ratio. For external division you make 1 part in the ratio negative... it doesn't matter which part.